At 1 July 2017, Lobstar Ltd acquired the following non-current assets:

Equipment $100 000 Vehicles $80 000

They are in different classes of non-current assets and are to be measured at fair value. The expected useful lives of vehicles and equipment are 5 years and 10 years, respectively.

At 30 June 2015, the fair values of both assets were assessed. The equipment had a fair value of $82 000, and the vehicles, $70 000. The remaining useful lives were assessed to be 8 years for equipment and 7 years for vehicles.

Required
Prepare the journal entries for Lobstar Ltd for the years ending 30 June 2015 and 2016.
Answer #1
Lobstar Ltd.
Date Particular Ref. Post Debit ($) Credit ($)
30-Jun-15 Depreciation Expense a/c                                                            Dr. W.N. 1 10,000
To Accumulated Depriciation a/c 10,000
(Depreciation charged on equipment)
Accumulated Depreciation a/c                                                  Dr. 10,000
To Equipment a/c 10,000
(Value of equipment written down to $90,000)
Loss on revaluation of equipment a/c                                    Dr. 8,000
To Equipment a/c 8,000
(Value of equipment decreased from $90,000 to $82,000)
Depreciation expense a/c                                                           Dr. W.N. 2 16,000
To Accumulated depriciation a/c 16,000
(Depreciation charged on vehicle)
Accumulated Depreciation a/c                                                  Dr. 16,000
To Vehicle a/c 16,000
(Value of equipment written down to $90,000)
Vehicle a/c                                                                                         Dr. 6,000
To gain on revaluation of Vehicle a/c                                     6,000
(Value of equipment increased from $64,000 to $70,000)
Income tax expense 1,800

Depreciation = Value of asset - Scrap Value / Useful life

W.N. 1: Depreciation on Equipment = 1,00,000 - 0 / 10 = $10,000

W.N. 2: Depreciation on Vehicle = 80,000 - 0 / 5 = $16,000

(P.S. Kindly give a thumbs up if you found the solution helpful.)